Home‎ > ‎Reports‎ > ‎Crew100B‎ > ‎

day05


Crew 100B Log Book for March 2 2011

Photos of the day





Commander's Check-In Report

Balwant Rai Reporting

Physical Status: Excellent

Experiment Status: All the medical, astrobiological, rover and food experiments began today with unique results.

Time Departed/Returned from EVA: four EVAs performed

The meals today were just filling since it was a non-cooking day.

EVA’s: We completed four successful EVA’s.

Number Of reports: 3 science reports, 1 engineering report, 1 journalist report, 4 EVA reports and 1 chef report. Plan for tomorrow: will initiate the medical experiment again on all crew members. Interpretation of the Sun by means of the radio telescope will also commence. The study of the rover’s maneuverability as well as the potential use of the drone, weather pending, will take place. Soil samples will yet again be collected during three EVA’s tomorrow. We discovered a dinosaur’s dig and will visit the same area with full preparation to collect useful evidences. One will be on foot, and two will use the ATV’s. Tomorrow will be a cooking day.


Commander's Report


Balwant Rai Reporting

Today we began our day with our continuing medical experiments as well as a new mission to locate fossils.

The crew’s medical health will continue to be monitored by Jasdeep for the remainder of our stay here on Mars. Breakfast was very agreeable and satisfying. Afterwards, some of the crew started the first EVA of the day with creating an obstacle course for the rover, and continuing to get familiar with flying the drone. The rest of our crew had their saliva samples taken (via Versi Sal 1 from Oasis Diagnostics), vital parameters monitored (via the Zephyr Bioharness), and their heart rate variability measured (via NERV Express 4.2 software). I also performed the cognitive study (via CogState Research software) on all crew members, as well.

Our crew scientists continued to work through problems with starting observations of sun via the radio telescope. Three more EVA’s came about, where soil samples were taken for the microbiology experiment. Later in the evening one more EVA occurred for a few of our crew members to explore another area of the Martian landscape. Today is a non-cooking day, so I am sure our chief and associate chef is busy finding something that might taste well. Tomorrow, we will continue our medical and astrobiological experiments, as well as continue working on getting the rover working well, and potentially getting the radio telescope up and working.

Engineering Report

Matthieu Ansart Reporting

Generator/Electricity:

· Kitty (Diesel Generator) (Used? Yes/No): Yes

· Honey (Gas Generator) (Used? Yes/No): No

· Battery Charge Level: Checked by DG

o 10.59 V everstart

o 12.37 V armor plate

o 13.32 V green

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

1 generator run, all is good


Fuel Status:

· Diesel (Full/0.75/0.5/0.25/Empty): 0.73

· Propane Tank (Full/0.75/0.5/0.25/Empty): 0.39

· Gasoline Tank (Full/0.75/0.5/0.25/Empty): 1 container full


Notes/Comments:


Oil:

· Kitty Oil Quantity (# of Quarts in storage): Leave it to DG

· Honey Oil Quantity (# of Quarts in storage): Leave it to DG

· ATV Oil Quantity (# of Quarts in storage): 6 quarts on the engineering work bench

· Car Oil Quantity (# of Quarts in storage):


Water Status:

· Outside Potable Water Tank Level (inches from bottom): 17

· Trailer Potable Water Tank Level (inches from bottom): 1/3 Full

· Hab Potable Water Tank Level (inches from bottom): 9

· Potable Water Meter Reading:


Notes/Comments: the problem with the toilets seems to be solved. We can flush it normally.


GreenHab:

· Condition of plants in Tank 1: green water

· Condition of plants in Tank 2: green things

· Condition of plants in Tank 3: green things

· Greenhab notes/ comments: Tank 1 is full. Some drops go into Tank 2


Observatory:

· Telescope Used? (Yes/no) No, observatory checked

Observatory Notes:


Transportation:

Hab Car (NOT RENTAL)

· Used (yes/no): No

· Oil Checked (yes/no): No

· Tire Status: Not checked

· Odometer at end of day: 123973

· Notes/Comments on Hab Car:


ATV 1 (Opportunity):

· Used: Yes

· Oil Checked: No

· Fuel Consumed:Replaced

· Tire Status: Filled

· Comments:


ATV 2 (Spirit):

· Used: No

· Oil Checked: No

· Fuel Consumed:

· Tire Status:

· Comments: Still not fixed


ATV 3 (Viking I):

· Used: Yes

· Oil Checked: No

· Fuel Consumed: Replaced

· Tire Status: Filled

· Comments:


ATV 4 (Viking II):

· Used: Yes

· Oil Checked: No

· Fuel Consumed:Replaced

· Tire Status: Filled

· Comments:


Heating and Ventilation:

· Thermostat setting upstairs (Farenheit): 65

· Actual temperature upstairs (Farenheit): 72

· Thermostat setting downstairs (Farenheit): 55

· Actual temperature downstairs (Farenheit): 63


Computers/Networking Infrastructure:

Notes/comments on internet/computers?:


General Engineering

1. Notes/Comments:


Science Report

Rachel Dompnier Reporting

Today I went for my first experiment on rover with Matthieu. This experiment was supposed to be the first test. Matthieu symbolized the novice driver. He had to follow a course with targets and avoid obstacles. I had to observe his reactions as a beginner facing difficulties.

Although he had never drive the rover before, Matthieu did not take so much time to get used to it. The first target was reached soon with some problems which occurred because of the delay of the cameras. Indeed, the rover hurt bushes, whereas Matthieu believed there were not any.

Besides, due to the low power of the rover, it has difficulties to go up even just small amounts of mud. The rover has to have a large amount of momentum to drive up some obstacles.

Although those problems that occurred, we are satisfied with this experiment widely successful. It enabled us to show which advantages the UAV could bring to the rover mission: it could help the rover to oversee obstacles and choose the best way to reach the target.

While Matthieu and I were busy with the rover, Quentin was practicing with the drone. He found a way to trim it and made adjustments which really improved the stability of the drone.


Science Report (Medical)

Jasdeep Kaur Reporting

Time: 1600-1730 : Blood pressure and pulse rate monitored by Zephyr technology and by manually inside Hab and after returning from EVA.

Crew Members: Science Team: 3 crew members Balwant Rai, Quentin Bourges and Jasdeep Kaur

Objective: To monitor the blood pressure by different exercises and massages while performing some pedestrian EVA to control vital body parameters.

Observation: Crew members were monitored for their blood pressures before going out for EVA. The blood pressures and heart rate was also monitored by Zephyr technology. Pulse was checked manually and electronically. The same measurements were taken again after returning back from EVA.

Results: As with respect to the principles to be followed on Mars, crew members have to monitor their regular and vital body functions. This needs very good maintainance of parameters like blood pressure, heart rate and pulse rate. This can be balanced by some exercise and massages. This is fifth day and it was found that during pedestrian EVA, crew members went really far up the heights and when they were back, their diastolic pressure raised, but it was in normal limits.

Lessons learned: Some of other yoga exercises and massages to be performed before departing for pedestrian EVAs.


Science Report (Bio-Medical)

Balwant Rai and Jasdeep Kaur reporting

Effects of mental stress on autonomic cardiac modulation during EVA in Mars Analog environment

The space missions like Mars missions necessitate the humans to have adaptability of their body to all extrinsic factors and to adapt to systemic and complex environments which are beyond human capacity and acceptance. This hypothesis is tested in Mars Analog simulated environment to study the alterations in effects of cardiovascular neural response to standardize cognitive activity.

All crew members were selected for this study. The Heart variability was measured before and after EVA by Zephyr BioHarness using (via NERV Express 4.2 software) analysis and a cognitive status by (via CogState Research software). Results concluded that score of tests is a valid mental stress tool for Marsonauts in MDRS and other mars analogous environment.


EVA 12 Report

Jasdeep Kaur reporting

EVA # 12, 10:55-13:00

Crew Members: Rachel Dompnier, Jasdeep Kaur, Balwant Rai

Site Location: Confidential

Transit Mode: ATVs

Objectives: Geo-Mission: to locate dinosaur's dig to find some bones and fossils

Results: At 10.45am in hab we could find some of unknown ATVs ouside, we went out of Hab and found some archaeologists with ATVs and they were going to locate some dinosaur's dig and fossils, we immediately followed them quite far and located the place. We couldnot stay there long as we had no time to take all equipments along,so we will be going in same direction tomorrow in a planned way and hope to collect some important information and bones.

Lessons Learned: As sometimes, the way is quite muddy, so its must to take mouth masks while driving with ATVs helmets to stop inhaling lots of mud and sand.


EVA 13 Report

Quentin Bourges Reporting

EVA#13 09:45 am to 10:35 am

Crew members: Rachel Dompnier, Matthieu Ansart, Quentin Bourges

Location : Green Hab and near by

Transit mode: Foot

Objective: Give a try for a novice with the Rover on a small obstacle course and detect difficulties encountered. In the same time, keep practicing with the UAV on the same obstacle course.

Results: The Rover lacks mechanical power because we could notice it got jammed each time the slope was very few degrees up. We also have a delay problem with the video camera that we cannot fix. Concerning the UAV, the trim problem has been fixed in the parameters panel control. We still have some difficulties to maneuver it as commands are very sensitive.

Lessons learned: We should adapt the operating speed with the Rover as the video camera delay cannot be fixed.


EVA 14 Report

Matthieu Ansart Reporting

EVA #15, 2:30 to 3:45 pm

Crew Members: Quentin Bourges, Crystal Latham, Matthieu Ansart,

Site Location: N 38.40798° W 110.82311°

Altitude : 1426 m

Transit Mode: ATV

Objectives: to take some samples, determine radio range and improve area knowledge.

Results: we bring some samples, shiny samples like crystals but crisper than quartz. We got another waypoint where we couldn't reach the Hab by radio.

Lessons Learned: we discover a west natural limit for our exploration. A cliff blocked our progression. We could improve our area knowledge in going to the North in order to cross this obstacle.


EVA 15 Report

Rachel Dompnier Reporting

EVA #15, 4:00 to 5:30 pm

Crew Members: Jasdeep Kaur, Quentin Bourgès, Balwant Rai

Site Location: near the Hab

Transit Mode: Foot

Objectives: Take soil samples. Record active blood pressure.

Results: Diastolic pressure increases but in safe limits.

Lessons Learned: Take water when you go on walking EVA. Secure good shoes to climb up.


Chef's Report

Quentin Bourges Reports

1a) Today was

- a non cooking day

1b) Was today a special day celebrated at the main meal with special food or activity? If so, what was special about today and what food(s) were prepared to mark the celebration?

No

2) List the foods served at today's main meal (usually the evening meal), giving full names of each.

Alpine Air “Kung Fu Chicken”

3) List any main meal foods not finished by the crew (leftovers stored for later use or discarded)

10% of kung fu chicken

4) Recipes for anything more complicated than following package directions exactly, or rehydrating.

Add some classic yellow mustard in with the kung fu chicken, it gives both spicy and sweetie taste.

5) Lessons learned (if any)

No

6) Comments/questions for Kim and Jean, the food study investigators

No


Journalist Report

Crystal Latham Reporting

Mission day five on Mars, was quite exciting. Today marked the first successful rotation of one of our unique medical experiments to all crew members, that happens daily. Other medical experiments, like monitoring blood pressure while on a walking EVA and monitoring the cognitive state of all crew members, are going really well.

Today there were four very successful EVA’s, which resulted in interesting data for both our medical and astrobiological experiments. Many interesting soil and rock samples were taken from various locals that may have some interesting information on the history of Mars and the life that exists here. Speaking of life on Mars, today some aliens came for a visit and showed some of our crew to a unique fossil site, that we hope to explore tomorrow.

Today was an extremely busy day for some of the crew since a three abstracts were due to the International Astronautical Federation, by 13:00 Mars time. These abstracts are related to the many medical experiments we are doing here at MDRS. These experiments have some important partners, like the National Aeronautical Space Agency, the JBR Research Group, and the European Space Agency.

It looks as though crew 100B is going to have a very successful rotation here on Mars. Our chief scientist finally got the radio Jove telescope up and running. So over the remaining ten days of our mission she will be looking for solar storms. The crew has designated mission day five as leftover day, since it was a non-cooking day. The rest of our stay here at MDRS is looking to be research filled. Crew 100 B is missing all the earthlings who are reading this.